# Adaptively Optimizing the Performance of HPX’s Parallel Algorithms

### v1

@@ 4.2 Customization Points for HPX’s Parallel Algorithms
As outlined in Section 2, in order to optimize a parallel algorithm, two parameters need to be adaptively controlled: the size of the chunks, i.e. the size of a subsection of the data array that is passed to a single concurrent task, and the number of processing units. Here we assume that the amount of work in the user-supplied loop body is either known or can be measured during the first invocation of the parallel algorithm and this data can be used to compute the optimal parameters.
@@
HPX’s parallel algorithm implementations rely on invoking several customization points. We will focus, in particular, on three of those: measure_iteration, processing_units_count, and get_chunk_size (see [7]). Listing 1 shows a simplified call sequence taken from the algorithm implementation of HPX. All of the customization points are called by passing the execution parameters object params and the executor exec that were bound to the execution policy used for invoking the parallel algorithm (see Section 4.1). The variable loop_body is a function representing the user-supplied loop body, while count holds the overall number of iterations.
// Return the time per iteration for a given
// loop body and overall number of iterations.
@@
// Return number of cores to utilize for a given time
// per iteration and overall number of iterations.
@@
// Return the chunk size for the given time per iteration,
// processing unit count, and overall number of iterations
@@
Listing 1: Exemplar invocation sequence of customization points in the implementation of HPX’s parallel algorithms
@@
  • measure_iteration should return the time spent execute the loop body.
  • processing_units_count should return the number of processing units the algorithm implementation should utilize.
  • get_chunk_size should return the number of array elements per task.
The default implementations for these customization points splits the work into equally sized chunks while utilizing all available processing units.
In Section 5 we present the results of implementing a customization point object adaptive_core_chunk_size (acc) that exposes the aforementioned customization functions by encapsulating the mathematical model we describe in Section 3. We will show that such an automatic customization is possible without changing the existing APIs implemented by HPX which are fully conforming to the parallel algorithms as defined by the C++ Standard. The required functionality is encapsulated in a simple C++ type that serves as the customization point for HPX’s parallel algorithms, without having to modify the algorithm’s implementation.
@@ 5 Experiments and Methodology
Our new executor is adapting the optimal number of cores and chunk size, so we call it adaptive_core_chunk_size or (acc) in this paper. In this execution parameters object, we have incorporated a customization point—measure_iteration—designed to compute the time per loop execution (See Listing 1). The time ($T_{i})$ will be calculated once for each workload, and then will be used to find $T_{1}$ which is the total time for a given number of iteration. HPX runs a benchmark on an empty thread to calculate overhead which is $T_{0}$. Another customization point used in this executor is processing_units_count (listing  1) that uses $T_{1}$ and $T_{0}$ and applies Equation 7 to find the optimal number of cores for each workload. It then uses that value, unless it is more than the maximum available cores in the system, in which case the maximum available cores are used. The last critical customization point in this executor is get_chunk_size which calculates optimal chunk size based on the calculated cores in processing_units_count. This Customization point uses the minimum time per chunk calculated based on Eq 10 and makes sure to have at least that chunk size for the optimal result (see Section 4.2).
For those more familiar with OpenMP than with HPX, the above implementation would provide an automatic way to pick the best value for the num_threads argument to the parallel pragma. Not only will this avoid slowdowns when loops are too small or quick to benefit from parallelism, but it leaves cores available for other parallel tasks should they be needed.
In order to better understand the behavior of map-type algorithms, we conducted a thorough study of the effects of core count and chunk size on the adjacent-difference algorithm, which is memory-bound, and artificial-work which is compute bound.
@@
The rationale behind this is that for a smaller workload, an increased number of cores and chunks escalates overhead, hindering performance, while for larger workloads with a larger value of $N_{C}$, more chunks enables greater load balancing. To maintain optimal performance, we use an executor that automatically adjusts the optimal number of cores and chunks.
@@
We evaluate all the performance measurements in this experiment using HPX V1.10.0 on a test machines with Intel Xeon Skylake processors, with 40 cores at 2.4GHz and 96 Gb of main memory, 2 sockets with 20 cores each, with hyperthreading disabled. In these experiments, each processor unit is the same as one core or thread. We used a benchmark to generate data and the result is average of 50 iterations.
In the initial experiment, we kept the number of cores constant while varying the chunk sizes for various workloads. HPX has a very light-weight parallelism with very efficient work stealing [1]. As a result, we expect to see good resource utilization. However, caution is needed, as excessive chunking can introduce significant overhead. The graphs also indicate the best chunking options for each number of cores used. We tested each number of cores with three possible chunking configurations: a number of chunks equal to the number of cores ($C=1$), four times the number of cores ($C=4$), and eight times the number of cores ($C=8$). See Figure 1.
@@
In the second experiment, we compared the different fixed core counts using $C=4$ along with the result with our adaptive algorithm using adapted cores and $C=8$ (acc). See Figure 2. This is equivalent to using OpenMP and providing static scheduling and a num_threads argument.
@@
Figure 2: Speedup measured for the adjacent difference algorithm across a range of cores counts and input sizes. We compare executions (for different numbers of cores) with the results measured when using the new adaptive_core_chunk_size (acc) (red line).
Experiment 2
Experiment 2 used 2 different hardwares, one is the same as experiment one, and the other is AMD EPYC processors with 48 cores, 2 sockets with 24 cores each. However, to evaluate the impact of our adaptive_core_chunk_size executor on compute-bound loops, we used an artificial workload instead of adjacent-difference. We then compared the performance of our new executor against the default parallel execution policies (used by both OpenMP and HPX) across varying workloads.
Figure 4, and  4 present this comparison. In these graphs, parallelization begins earlier and scales more rapidly. As before, $T_{1}$ represents the total execution time. However, this algorithm scales differently compared to the previous case (bigger $T_{1}$ for the same input size). As a result, the adaptive_core_chunk_size execution parameters object not only starts parallelization on smaller arrays, but it also uses all available cores in the system earlier than in the default case. As a result, especially very small workloads show significant benefits from dynamic core adjustments with our new executor. Additionally, our executor consistently outperforms other core configurations in larger workloads. This performance gain is attributed to its ability to select optimal chunk sizes compared to the default parallel policies in HPX.
Figure 3: Speedup across various core counts for a compute-bound use case when using the default static parameters compared to using the new adaptive_core_chunk_size(acc)(red line) across varying input sizes on an Intel hardware.
Figure 4: Speedup across various core counts for a compute-bound use case when using the default static parameters compared to using the new adaptive_core_chunk_size(acc)(red line) across varying input sizes on an AMD hardware.
@@ 6 Results
Performance analysis of varying chunk sizes. In the first experiment, we varied chunk sizes while keeping the number of cores constant. The results, illustrated in Figures 1(a)1(b), and 1(c) highlight how performance varies with chunk sizes across various workloads. Having the chunks eight times of number of cores is always better option.
@@
Performance of adaptive chunking. Building on the insights from the two experiment sets, we developed an adaptive_core_chunk_size customization point object. This object dynamically adjusts core and chunk configurations based on workload size. Figure 2 illustrates the performance of various processor counts alongside the results of our executor. The red line (acc) optimally selects the number of cores and chunks for each workload. By dynamically adjusting the number of processors—utilizing fewer for smaller workloads and more for larger ones—this approach demonstrates improved overall performance across all workloads.
@@
It is also worth noting that in Experiment 1, figure 2, although our new executor achieves the best overall performance, the parallel efficiency remains limited, even with 40 cores. Specifically, we observe only approximately a 10x speedup compared to the sequential execution. This is primarily because adjacent_difference is a memory-bound algorithm. However, when the algorithm is changed to a compute-bound one, as shown in figures 4 and 4, the benefits of parallelism become more evident. On a 40-core machine, we achieve up to 38x speedup, and on a 48-core machine, the speedup reaches up to 46x compared to the sequential execution.
@@
In summary, the new approach is beneficial in three aspects: (1) performance gain, the adaptive executor consistently outperformed static configurations. By dynamically tuning core and chunk sizes, it minimized overheads and maximized resource utilization, leading to significant performance improvements across all workload sizes. (2) overhead reduction, for small workloads, the adaptive executor effectively reduced overhead by limiting cores and chunks. For large workloads, it leveraged maximum available cores and increased chunk count, enhancing parallelism. (3) Adaptability, the adaptive executor’s ability to adjust to varying workloads makes it highly versatile, suitable for a wide range of parallel applications.
@@ 7 Conclusions
The development and optimization of parallel algorithms are crucial for maximizing computational resource efficiency. This paper studies map-type algorithms and develops a mathematical model to improve their performance based on runtime parameters. The focus of the study is on determining the optimal combination of core counts and chunk sizes at runtime for various workloads. The model has been implemented and integrated in the HPX C++ library. By introducing the HPX execution parameters object adaptive_core_chunk_size, we have made it possible to dynamically adjust the number of cores and the size of the chunks based on the application’s workload size and the available number of cores in the execution environment. The result is improved performance and scalability across a wide range of workloads for parallel algorithms. This adaptive approach reduces the overheads and enhances resource utilization, leading to better overall performance. This is particularly true for smaller workloads.
@@ 8 Future Work
Future work will focus on several areas to build upon the findings of this paper. First, we aim to explore the application of the adaptive_core_chunk_size execution parameters object to other algorithms. This includes investigating its impact on additional memory-bound and I/O-bound algorithms to assess its versatility and performance benefits across different computational paradigms. Second, fluctuations in the performance graphs suggest the influence of cache effects. Future models can factor in these effects by leveraging measurement tools like PAPI.
Lastly, we plan to conduct comprehensive benchmarking on a variety of real-world applications to validate the executor’s effectiveness and robustness. This includes performance evaluations on different hardware architectures, such as GPUs and specialized accelerators, ensuring the executor’s adaptability and efficiency across diverse computing environments. By testing on a broader range of platforms, we aim to demonstrate the generalizability and scalability of our approach.

### v2

@@ 4.1.3 Customization Points for HPX’s Parallel Algorithms
As outlined in Section 2, in order to optimize the execution times for parallel algorithms, two parameters need to be adaptively controlled: the size of the chunks, i.e. the size of the parts of the data array the algorithm is invoked on that are passed as separate concurrent tasks to a processing unit, and the number of processing units to utilize for a particular execution of a parallel algorithm. In this work we assume that the amount of work in the user-supplied loop body can be measured during the first invocation of a particular, repeatedly invoked parallel construct.
@@
HPX’s parallel algorithm implementations rely on invoking several customization points. We will focus, in particular, on three of those: measure_iteration, processing_units_count, and get_chunk_size (see [14]). Listing 3 shows a simplified call sequence taken from the algorithm implementation of HPX. All of the customization points are called by passing the execution parameters object params and the executor exec that were bound to the execution policy used for invoking the parallel algorithm (see Section 4.1.1). This allows the programmer to provide the necessary execution context to the customization point implementations. The variable loop_body is a function representing the user supplied loop body, while count holds the overall number of iterations.
// Return the time per iteration for given
// loop body and overall number of iterations
@@
// Return number of cores to utilize for
// given time per iteration and overall
// number of iterations
@@
// Return size of chunks (i.e. number of
// iterations per chunk) to run as concurrent
// tasks for given time per iteration, number
// of processing units, and overall number of
// iterations
@@
Listing 3: Exemplar invocation sequence of customization points in the implementation of HPX’s parallel algorithms
@@
  • measure_iteration should return the time required to execute one iteration of the loop in some unspecified way.
  • processing_units_count should return the number of processing units the algorithm implementation should utilize while running the overall algorithm.
  • get_chunk_size should return the sizes of the array parts (chunks) that are run as concurrent tasks.
The default implementations for these customization points perform minimal work resulting in splitting the array into chunks such that each of the available processing units receives an equal amount of iterations.
ere we present the results of implementing an execution parameter type adaptive_core_chunk_size that exposes the aforementioned customization points by encapsulating the mathematical model we describe in Section 3.
@@ 5 Experiments and Methodology
Our new executor adaptively finds the optimal number of cores and the chunk size, so we call it adaptive_core_chunk_size or (acc) in this paper.
In this execution parameters object, we have implemented the customization point—measure_iteration—designed to compute the time per loop execution (See Listing 3). The value of $T_{1}$ will be calculated once for each workload, the first time a loop is invoked. (Note that because the number of elements, $N_{E}$, may vary with different invocations of the loop, what we actually store is $T_{E}=T_{1}/N_{E}$. In this way, if future invocations of the loop have a different $N_{E}$, we can easily calculate the appropriate $T_{1}$ for that invocation.)
$T_{0}$ is an application constant value and it should be measured on a particular application and particular architecture. When we run an application on a particular architecture, we measure $T_{0}$ on the first invocation loop and use this value to optimize future invocations. We make sure this value is measured each time for each application separately since this value can be different for different loops.
The processing_units_count customization point (listing 3) uses $T_{1}$ and $T_{0}$ and applies Equation 7 to find the optimal number of cores, $N$, for each workload. It then uses that value, unless it is more than the maximum available cores in the system, in which case the maximum available number of cores are used. The last critical customization point implemented by this executor is get_chunk_size, which calculates the optimal chunk size based on the calculated number of cores (as previously returned from processing_units_count). This Customization point uses the minimum time per chunk calculated based on Eq 9 and makes sure to return a chunk size ensuring the optimal result (see Section 4.1.3).
For those more familiar with OpenMP than with HPX, the above implementation would provide an automatic way to pick the best value for the num_threads argument to the parallel pragma. Not only will this avoid slowdowns when loops are too small to benefit from parallelism, but it leaves cores available for other parallel tasks should they be needed.
In order to better understand the behavior of map-type algorithms, we conducted a thorough study of the effects of core count and chunk size on the adjacent-difference algorithm, which is memory-bound, and an artificial-work algorithm that is compute bound and an algorithm in ChplX.
@@
The rationale behind this is that for a smaller workload, an increased number of cores and chunks escalates overhead, hindering performance, while for larger workloads with a larger value of $N$, more chunks enables greater load balancing. To maintain optimal performance, we use an executor that automatically adjusts the optimal number of cores and chunks.
@@
We evaluated all the performance measurements in this experiment using HPX V1.10.0 on a test machine with Intel Xeon Skylake processors, with 40 cores at 2.4GHz and 96 Gb of main memory, 2 sockets with 20 cores each, with hyperthreading disabled. In these experiments, each processor unit is the same as one core or thread. We used a benchmark to generate data and the result is the average of 50 iterations.
In the initial experiment, we kept the number of cores constant while varying the chunk sizes for various workloads. HPX has a very light-weight parallelism with very efficient work stealing [15]. As a result, we expect to see good resource utilization. However, caution is needed, as excessive chunking can introduce significant overhead. The graphs also indicate the best chunking options for each number of cores used.
In our first paper [4], we evaluated core counts of 2, 16, and 32 with three possible chunking configurations: a number of chunks equal to the number of cores ($C=1$), four times the number of cores ($C=4$), and eight times the number of cores ($C=8$). See Figure 1(a), 1(b), and 1(c). In this paper, we extended our evaluation to utilize all available cores on the test machine (see Figure 1(d)). The results indicate that, among the fixed chunk configurations, $C=8$ consistently yielded the best overall performance. We also explored finer-grained chunking by comparing the optimal setting of $C=8$ against larger values—specifically, 16, 32, and 64 times the number of cores (C=16,32,64C=16,32,64)—as shown in Figure 1(e).
@@
(d) Speedup with 40 processing units (cores) (New)
(e) Speedup with 40 processing units (cores) (New)
@@
Experiment 2
In the second experiment, we compared the different fixed core counts using $C=4$ (current HPX implementation) along with the result with our adaptive algorithm using adapted cores and $C=8$ which was the most optimal chunking (acc). See Figure 2. This is equivalent to using OpenMP and providing static scheduling and a num_threads argument. All the performance measurements in this experiment are the same as Experiment 1.
@@
Figure 2: Speedup measured for the adjacent difference algorithm across a range of core counts and input sizes. We compare executions (for different numbers of cores) with the results measured when using the new adaptive_core_chunk_size (acc) (red line).
Experiment 3
In our initial result [4], we ran this experiment on two different machines. In this paper we use three different types of hardware: one is the same as experiment one (Intel Hardware); the second is AMD EPYC processors with 48 cores, 2 sockets with 24 cores each; and the third (the new one) is a RISC-V architecture with 64 cores and four Numa nodes. However, to evaluate the impact of our adaptive_core_chunk_size executor on compute-bound loops, we used an artificial workload instead of adjacent-difference. We then compared the performance of our new executor against the default parallel execution policies (used by both OpenMP and HPX) across varying workloads.
Figure 5, 5 , and 5 present this comparison. In these graphs, parallelization begins earlier and scales more rapidly. As before, $T_{1}$ represents the total execution time. However, this algorithm scales differently compared to the previous case (bigger $T_{1}$ for the same input size). As a result, the adaptive_core_chunk_size execution parameters object not only starts parallelization on smaller arrays, but it also uses all available cores in the system earlier than in the default case. As a result, especially very small workloads show significant benefits from dynamic core adjustments with our new executor. Additionally, our executor consistently outperforms other core configurations in larger workloads. This performance gain is attributed to its ability to select optimal chunk sizes compared to the default parallel policies in HPX.
Figure 5 presents our new experiment conducted on a RISC‑V machine. We observed that the time per iteration on this platform is noticeably higher than on our other test machines, which may be attributed to specific architectural constraints of RISC‑V systems [21]. Regarding this characteristic—namely, that RISC-V is inferior in computation time compared to platforms like AMD [22]—our executor still outperforms other experimental cases.
Figure 3: Speedup across various core counts for a compute-bound use case when using the default static parameters compared to using the new adaptive_core_chunk_size(acc)(red line) across varying input sizes on Intel hardware.
Figure 4: Speedup across various core counts for a compute-bound use case when using the default static parameters compared to using the new adaptive_core_chunk_size(acc)(red line) across varying input sizes on AMD hardware.
Figure 5: Speedup across various core counts for a compute-bound use case when using the default static parameters compared to using the new adaptive_core_chunk_size(acc)(red line) across varying input sizes on RISC-V hardware. (New)
Experiment 4
In this new set of experiments, we used the ChplX library, which is an abstraction that mimics chapel’s semantics [5]. It is designed for a line by line source to source translation of Chapel programs to C++ with HPX as its runtime system. More about this is explained in [5] paper. In this library, we chose ChplX forall as an algorithm for our new experiment. Figure 6 shows the comparison between our new executor acc with different number of cores and static number of chunks used in HPX. The machine used for this experiment is AMD as defined earlier.
Figure 6: Speedup measured for the Chplx forall across a range of core counts and input sizes. We compare executions (for different numbers of cores) with the results measured when using the new adaptive_core_chunk_size (acc) (red line). (New)
@@ 6 Results
Performance analysis of varying chunk sizes. In the first experiment, we varied chunk sizes while keeping the number of cores constant. The results, illustrated in Figures 1(a)1(b)1(c), 1(d), and 1(e) highlight how performance varies with chunk sizes across various workloads. Setting the number of chunks to eight times of number of cores is always the better option.
(a) Strong Scaling for different core counts vs acc Executor
(b) Strong Scaling for different core counts vs acc Executor
(c) Strong Scaling for different core counts vs acc Executor)
Figure 7: Strong scaling results for three input sizes. Speedup is plotted against core count (including acc) for fixed input sizes labeled in $log_{2}$ scale. The sequential baseline is normalized to 1. (New)
@@
Performance of adaptive chunking. Building on the insights from the two experiment sets, we developed a customization point object adaptive_core_chunk_size. This object dynamically adjusts core and chunk configurations based on workload size. Figure 2 illustrates the performance of various processor counts alongside the results of our executor. The red line (acc) optimally selects the number of cores and chunks for each workload. By dynamically adjusting the number of processors—utilizing fewer for smaller workloads and more for larger ones—this approach demonstrates improved overall performance across all workloads.
@@
It is also worth noting that in Experiment 2, Figure 2, although our new executor achieves the best overall performance, the parallel speedup remains limited, even with 40 cores. Specifically, we observe only approximately a 10x speedup compared to the sequential execution. This is primarily because the algorithm adjacent_difference is a memory-bound algorithm and there is not enough work to maintain 95% efficiency. However, when the algorithm is changed to a compute-bound one, as shown in Figures 5, 5, and 5 the benefits of parallelism become more evident. On a 40-core machine, we achieve up to 38x speedup, on a 48-core machine, the speedup reaches up to 46x, and on a 64-core machine, we achieve up to 57x speedup compared to the sequential execution.
In our new experiments, We also used strong scaling as a tool to illustrate how core count affects parallel speedup and to compare the performance of our new executor, acc, against standard configurations. See Figure 7. We selected three input sizes—ranging from small to large—to represent different workload scales. The strong scaling results demonstrate that for smaller input sizes, using a large number of cores is suboptimal, whereas larger inputs benefit from increased core counts. Our adaptive executor, acc, consistently achieved the best speedup by adjusting the core count accordingly: it selected fewer cores for smaller workloads and more cores for larger ones. This adaptive behavior highlights acc’s ability to match resource usage to workload characteristics, maximizing performance across varying scenarios.
@@
In summary, the new approach is beneficial in three key ways. Firstly, through performance gain, the adaptive executor consistently outperformed static configurations. By dynamically tuning core and chunk sizes, it minimized overheads and maximized resource utilization, leading to significant performance improvements across all workload sizes. Secondly, through overhead reduction, for small workloads, the adaptive executor effectively reduced overhead by limiting cores and chunks. For large workloads, it leveraged maximum available cores and increased chunk count, enhancing parallelism. Thirdly, through daptability, the adaptive executor’s ability to adjust to varying workloads makes it highly versatile, suitable for a wide range of parallel applications.
@@ 7 Conclusions
The development and optimization of parallel algorithms are crucial for maximizing computational resource efficiency. This paper studies map-type algorithms and develops a mathematical model to improve their performance based on runtime parameters. The focus of the study is on determining the optimal combination of core counts and chunk sizes at runtime for various workloads. The model has been implemented and integrated into the HPX C++ library. By introducing the HPX execution parameters object adaptive_core_chunk_size, we have made it possible to dynamically adjust the number of cores and the size of the chunks based on the application’s workload size and the available number of cores in the execution environment. The result is improved performance and scalability across a wide range of workloads for parallel algorithms. This adaptive approach reduces the overheads and enhances resource utilization, leading to better overall performance. This is particularly true for smaller workloads.
@@ 8 Future Work
Future work will focus on several areas to build upon the findings of this paper. First, we aim to explore the application of the adaptive_core_chunk_size execution parameters object to more algorithms. This includes investigating its impact on additional memory-bound and I/O-bound algorithms to assess its versatility and performance benefits across different computational paradigms. Second, fluctuations in the performance graphs suggest the influence of cache effects. Future models can factor in these effects by leveraging measurement tools like PAPI. In this work, we develop an abstract model optimizing the performance of parallel loop construct in general cases. We also showed the models applicability for a set of artificial use cases. Applying this model to a real world application is part of our planned future work and is thus beyond the scope of this paper. We plan to conduct comprehensive benchmarking on a variety of real-world applications to validate the executor’s effectiveness and robustness. This includes performance evaluations on different hardware architectures, such as GPUs and specialized accelerators, ensuring the executor’s adaptability and efficiency across diverse computing environments. By testing on a broader range of platforms, we aim to demonstrate the generalizability and scalability of our approach. We also plan to integrate machine learning techniques to enhance the executor’s decision-making process. By training models on various workloads and system configurations, the executor can predict and adapt to optimal core and chunk sizes more accurately and dynamically. This approach could further reduce overheads and improve performance, particularly in heterogeneous computing environments. Moreover, additional experiments across different architectures, algorithms, and libraries revealed that while acc consistently performs well over a range of input sizes, there are still cases where it could make more optimal decisions regarding core selection. We plan to further investigate and refine the underlying formula to improve its performance across all input sizes.